OpenStudy (elenathehomeschooler):

Use the half-angle formula to evaluate tan(17pi/12)

OpenStudy (mathmale):

Please look up, and then share here, the formula for the tangent of a half angle. You could look up "half angle tangent formula." Once you have that formula, we'll proceed to use it to find tan (17pi / 12).

OpenStudy (elenathehomeschooler):

\[\tan a/2 = \pm \sqrt{1-\cos a/1+\cos a}\]

OpenStudy (elenathehomeschooler):

@mathmale

OpenStudy (jtug6):

As reference, you could also use \[(1-\cos(a))/(\sin (a))\] as well, as they're both the same formula however this one is slightly easier =)

OpenStudy (elenathehomeschooler):

can you help me @jtug6

OpenStudy (jtug6):

Uhhh sure! I generally don't like to take away from people who responded here first but since you asked =). Okay so we know that the half-angle identity is (1-cos(a))/(sin(a)) right? And we were given tan(17pi/12) as well, right?

OpenStudy (elenathehomeschooler):

yes.

OpenStudy (jtug6):

Okay. Well we need to account for the fact that we have a/2 and we're trying to get our tan(17pi/12) to be in that a/2 form, so we need to double our 17pi/12 and then divide that by 2. So i'll draw that real quick

OpenStudy (jtug6):